# Gravitational Waves in a Perturbative Bianchi I Background

### v1

@@ Abstract
It is straightforward to take the gravitational wave solution to first order in $v/c$ far from a binary source in a Minkowski background and adapt it to the Friedmann-Lemaitre-Robertson-Walker (FLRW) background, representing an expanding isotropic homogeneous universe. We find the analogous solution for a slightly anisotropic background, which may be a more accurate description of our universe, and we use a perturbative form of the Bianchi I metric and demonstrate how the waveform differs. Using supernova anisotropy data as a reference, we show that the assumption of a Bianchi I background could imply a 3.2% difference in inferred luminosity distance compared to what would be inferred under the assumption of the FLRW background. Therefore, the background spacetime used for the inference of the Hubble parameter from gravitational wave data should be considered carefully.
@@ I Introduction
The advent of the detection of gravitational waves by LIGO has brought incredible new opportunities for our understanding of the cosmos. It is well known that there is currently a Hubble tension, which is the disagreement of the estimate of the rate of the universe’s expansion inferred from the cosmic microwave background (CMB) in the early universe and supernovae in the late, local universe. The hope is that the determination of the Hubble parameter from gravitational wave (GW) data will give a more reliable value since it would be independent of the cosmic distance ladder, upon which supernovae analyses rely. Therefore, it is important to infer the luminosity distance of GW sources accurately as they determine the inferred Hubble parameter. A multiplicity of gravitational wave detectionsGhosh ; Pierra ; LIGO2017 and multimessenger detectionsMelo ; LIGOmm will afford a precise measurement of the Hubble parameter that is independent of the somewhat shaky rungs of the cosmic distance ladderCamarena .
@@
Obtaining the Hubble parameter from the CMB, supernovae, and GWs all assume the Friedmann-Lemaitre-Robertson-Walker (FLRW) metric and general relativity (GR) conventionally. However, there have been careful statistical analyses of supernovae, quasar, and CMB data that suggest that an isotropic and homogeneous FLRW universe is not accurate to as high a degree as is usually assumed Secrest:2022uvx ; Mohayaee:2021jzi ; Secrest:2020has ; Colin:2019opb ; Cea:2022zep ; Krishnan:2021jmh ; Lombriser:2019ahl . For example, in Colin:2019opb , the luminosity distance is expressed in a kinematic Taylor series to avoid making assumptions about the matter content or dynamics, and the inferred acceleration is essentially in the direction of the local peculiar bulk flow. Other works that argue for negligible evidence for anisotropy Brout:2022vxf ; Akarsu:2019pwn typically assume an isotropic framework to some degree or make corrections for peculiar velocities by transforming to the “CMB frame”, which removes the evidence for anisotropy. In Secrest:2020has , the authors show that the distribution of distant matter as traced by radio sources and quasars is not isotropic in the CMB frame. The authors of Secrest:2022uvx demonstrate a $5.1\sigma$ rejection of the null hypothesis that there is a dipole anisotropy in the distribution of radio galaxies and quasars consistent with our inferred peculiar motion. The question of bulk anisotropy may not be settled in the literature, but at the very least, metrics that account for anisotropy should be considered and studied.
@@
The accuracy of predictions of the waveforms of GWs due to massive binary systems has been refined to a high level over the years through advancements in numerical relativity. In the duration between emission of these waves from the source and detection by a gravitational wave (GW) observatory, the waveform is to a large degree unaffected by the spacetime through which it propagates. However, as more detections of GWs are made and as more observatories come online with higher precision, our ability to discern subtle effects on the waveforms will increase. In this work, using supernova anisotropy data as a reference, we show that the assumption of a Bianchi I background spacetime could imply a 3.2% difference in inferred luminosity distance compared to what would be inferred under the assumption of the FLRW background. The effect on the waveform is primarily in the amplitude as opposed to a more subtle effect on the oscillation pattern.
@@ II Bianchi I Metric
The simplest possible expanding universe that obeys the cosmological principle (homogeneous, isotropic) is described by the the FLRW metric,
@@
where $a$, $b$ and $c$ are arbitrary functions of time. We have chosen to work with a perturbative form of this metric for the ease of comparing the waveform to the solution with an FLRW background, where $b$ and $c$ differ from $a$ only by small constants:
@@
$c(t)=a(t)(1+\gamma).$
We assume that $\beta$ and $\gamma$ are also slowly varying. Therefore, $\beta$ and $\gamma$ satisfy
$\dot{\beta}\approx\dot{\gamma}\approx\beta^{2}\approx\gamma^{2}\approx 0.$ (4)
@@ II.1 Gravitational Waves
Without loss of generality, we consider a wave propagating in the $x$ direction. We assume the transverse-traceless (TT) gauge. To first order in $h_{\mu\nu}$ and next-to-leading order in wavelength compared to the curvature length scale, the solution to the wave equation has two non-zero independent components, $h_{22}$ and $h_{23}$, given by Eqs. (55) and (56) from Mondal:2020iet :
$\square h_{22}+2\left(\frac{b\dot{a}\dot{b}}{b^{3}}h_{11}+\frac{b\dot{b}\dot{c}}{c^{3}}h_{33}-b\ddot{b}~{}h_{00}\right)-\frac{2}{b}\left(\ddot{b}+\frac{\dot{a}\dot{b}}{a}+\frac{\dot{b}\dot{c}}{c}\right)h_{22}=0,$ (6)
$\square h_{23}-2\frac{\dot{b}\dot{c}}{bc}h_{32}-\frac{1}{b}\left(\ddot{b}+\frac{\dot{a}\dot{b}}{a}+\frac{\dot{b}\dot{c}}{c}\right)h_{23}-\frac{1}{c}\left(\ddot{c}+\frac{\dot{a}\dot{c}}{a}+\frac{\dot{b}\dot{c}}{b}\right)h_{32}=0.$ (7)
The propagation in the $x$ direction, the transverse condition, the traceless condition, and the symmetry of the metric tensor imply the following conditions:
@@
These conditions allow Eqs. (6) and (7) to be expressed as
$\left(\square-2\frac{b\dot{b}\dot{c}}{c^{3}}-\frac{2}{b}(\ddot{b}+\frac{\dot{a}\dot{b}}{a}+\frac{\dot{b}\dot{c}}{c})\right)h_{22}=0,$ (9)
$\left(\square-2\frac{\dot{b}\dot{c}}{bc}-\frac{1}{b}(\ddot{b}+\frac{\dot{a}\dot{b}}{a}+\frac{\dot{b}\dot{c}}{c})-\frac{1}{c}(\ddot{c}+\frac{\dot{a}\dot{c}}{a}+\frac{\dot{b}\dot{c}}{b})\right)h_{32}=0.$ (10)
Under the local wave approximation, the background dynamics are slow compared to the angular frequency $\omega$ of the wave:
$\mathcal{O}(\frac{1}{t^{2}}h_{\mu\nu})\ll\mathcal{O}(\omega^{2}h_{\mu\nu}).$ (11)
The time derivative of any of the scale factors goes like $1/t$, and so the second term of Eq. (9) goes like $h_{22}/t^{3}$, and the third term of Eq. (9) and all but the first term of Eq. (10) go like $h_{\mu\nu}/t^{2}$. These are all negligible compared to the d’Alembertian, since $\square h_{\mu\nu}\sim\omega^{2}h_{\mu\nu}$. Therefore, the wave equation simplifies to
$\square h_{22}\approx 0,~{}~{}\square h_{32}\approx 0.$ (12)
Let $\phi$ below stand in for $h_{22}$ and $h_{32}$. The D’alembertian operator unpacks to
$\square\phi=(\frac{1}{a^{2}}\frac{d^{2}}{dx^{2}}+\frac{1}{b^{2}}\frac{d^{2}}{dy^{2}}+\frac{1}{c^{2}}\frac{d^{2}}{dz^{2}})\phi-(\frac{\dot{a}}{a}+\frac{\dot{b}}{b}+\frac{\dot{c}}{c})\dot{\phi}-\ddot{\phi}.$ (13)
Using Eqs. (3) and (4), truncating to first order in $\beta$ and $\gamma$, we get
$\square\phi\approx\frac{1}{a^{2}}(\frac{d^{2}}{dx^{2}}+(1-2\beta)\frac{d^{2}}{dy^{2}}+(1-2\gamma)\frac{d^{2}}{dz^{2}})\phi-3\frac{\dot{a}}{a}\dot{\phi}-\ddot{\phi}.$ (14)
@@
$\square\phi=\frac{1}{a^{2}}(\frac{d^{2}}{dx^{2}}+(1-2\beta)\frac{d^{2}}{dy^{2}}+(1-2\gamma)\frac{d^{2}}{dz^{2}})\phi-3\frac{a^{\prime}}{a^{3}}\phi^{\prime}+\frac{a^{\prime}}{a^{3}}\phi^{\prime}-\frac{1}{a^{2}}\phi^{\prime\prime}.$ (16)
We write this as
$\square\phi=\frac{1}{a^{2}}\tilde{\nabla}^{2}\phi-2\frac{a^{\prime}}{a^{3}}\phi^{\prime}-\frac{1}{a^{2}}\phi^{\prime\prime},$ (17)
@@
~$\tilde{\nabla}^{2}\equiv\frac{d^{2}}{dx^{2}}+(1-2\beta)\frac{d^{2}}{dy^{2}}+(1-2\gamma)\frac{d^{2}}{dz^{2}}.$ (18)
The equation to solve is now
~$\tilde{\nabla}^{2}\phi-2\frac{a^{\prime}}{a}\phi^{\prime}-\phi^{\prime\prime}=0.$ (19)
@@
$\phi(x,y,z,\eta)=\frac{g(\eta-\tilde{r})}{\tilde{r}a(\eta)},$ (20)
@@
This $\tilde{r}$ is very similar to the Cartesian $r$ in the FLRW background. The factors of $\beta$ and $\gamma$ which scale the scaling factor $a$ for the $y$ and $z$ directions can be thought of equivalently as rescalings of $y$ and $z$:
@@
~$\tilde{\nabla}^{2}\left(\frac{g}{\tilde{r}a}\right)+\frac{a^{\prime\prime}}{a}\frac{g}{\tilde{r}}-\frac{g^{\prime\prime}}{\tilde{r}}=0.$ (23)
The second term goes like
$\frac{a^{\prime\prime}}{a}\frac{g}{\tilde{r}}\sim\frac{1}{\eta^{2}}\frac{g}{\tilde{r}},$ (24)
which is very small compared to the other terms since the background expansion rate is much slower than the oscillation rate of the wave (see Eq. (11)). We now have the final form of the wave equation to solve, having applied the local wave approximation:
~$\tilde{\nabla}^{2}\left(\frac{g}{\tilde{r}a}\right)-\frac{g^{\prime\prime}}{\tilde{r}}=0.$ (25)
This is satisfied by our assumed form Eq. (20) to first order in $\beta$ and $\gamma$. The solution for the corresponding GW equation that results when an FLRW background is used (see Eq. 4.182 of Maggiore ) is the same as the solution Eq. (20) when
@@
The solution for a binary system in circular orbit to first order in $v/c$ can be made to include back-reaction with respect to an FLRW background in a simple way using the local wave approximation (Eqs. 4.191 and 4.192 from Maggiore ). Because we have the same solution if we substitute $\tilde{r}$ for $r$, we can also express the solution for the two polarizations of GWs for our perturbative Bianchi I (PBI) background:
$h_{PBI,+}=\frac{4}{a(t_{0})\tilde{r}}(1+\mathrm{z})^{2/3}(GM_{c})^{5/24}(\frac{5}{256}\frac{1}{\tau_{obs}})^{1/4}\cos{(-2(5GM_{c})^{-5/8}\tau_{obs}^{5/8}+\Phi_{0})}\frac{1+\cos^{2}\iota}{2},$ (27)
$h_{PBI,\times}=\frac{4}{a(t_{0})\tilde{r}}(1+\mathrm{z})^{2/3}(GM_{c})^{5/24}(\frac{5}{256}\frac{1}{\tau_{obs}})^{1/4}\sin{(-2(5GM_{c})^{-5/8}\tau_{obs}^{5/8}+\Phi_{0})}\cos\iota.$ (28)
The luminosity distance in FLRW space is
$d_{FLRW}=a(t)r,$ (29)
@@
$h_{FLRW,+}=\frac{4}{d_{FLRW}}(1+\mathrm{z})^{2/3}(GM_{c})^{5/24}(\frac{5}{256}\frac{1}{\tau_{obs}})^{1/4}\cos{(-2(5GM_{c})^{-5/8}\tau_{obs}^{5/8}+\Phi_{0})}\frac{1+\cos^{2}\iota}{2},$ (30)
$h_{FLRW,\times}=\frac{4}{d_{FLRW}}(1+\mathrm{z})^{2/3}(GM_{c})^{5/24}(\frac{5}{256}\frac{1}{\tau_{obs}})^{1/4}\sin{(-2(5GM_{c})^{-5/8}\tau_{obs}^{5/8}+\Phi_{0})}\cos\iota.$ (31)

### v2

# Inferred Hubble Parameter from Gravitational Waves in a Perturbative Bianchi I Background

@@ Abstract
It is straightforward to take the gravitational wave solution to first order in $v/c$ far from a binary source in a Minkowski background and adapt it to the Friedmann-Lemaitre-Robertson-Walker (FLRW) background, representing an expanding isotropic and homogeneous universe. We find the analogous solution for a slightly anisotropic background, which may be a more accurate description of our late universe through which gravitational waves propagate, and implications from tight CMB anisotropy constraints may not necessarily determine the level of anisotropy in the late universe in light of modified gravity models as well as the Hubble tension. We use a perturbative form of the Bianchi I metric and demonstrate how the waveform differs. Using supernova anisotropy data as a reference, we show that the assumption of a Bianchi I background could imply on average a 2.1% difference in inferred luminosity distance compared to what would be inferred under the assumption of the FLRW background. This difference can be as high as 5.9% depending on the observation direction. Therefore, the background spacetime used for the inference of the Hubble parameter from gravitational wave data should be considered carefully.
@@ I Introduction
The advent of the detection of gravitational waves by LIGO has brought incredible new opportunities to our understanding of the cosmos. It is well known that there is currently a Hubble tension, which is the disagreement of the estimate of the rate of the universe’s expansion inferred from the cosmic microwave background (CMB) in the early universe and supernovae in the late, local universe. The determination of the Hubble parameter from gravitational wave (GW) data should give a more reliable value since it would be independent of the cosmic distance ladder, upon which supernovae analyses rely. Therefore, it is important to infer the luminosity distance of GW sources accurately as they determine the inferred Hubble parameter. A multiplicity of gravitational wave detectionsGhosh ; Pierra ; LIGO2017 and multi-messenger detectionsMelo ; LIGOmm will afford a precise measurement of the Hubble parameter that is independent of the somewhat shaky rungs of the cosmic distance ladderCamarena .
@@
Obtaining the Hubble parameter from the CMB, supernovae, and GWs all assume the Friedmann-Lemaitre-Robertson-Walker (FLRW) metric and general relativity (GR) conventionally. However, careful statistical analyses of supernova, quasar, and CMB data have suggested that an isotropic and homogeneous FLRW universe is not as accurate to a degree as usually assumed Secrest:2022uvx ; Mohayaee:2021jzi ; Secrest:2020has ; Colin:2019opb ; Cea:2022zep ; Krishnan:2021jmh ; Lombriser:2019ahl . For example, in Colin:2019opb , the luminosity distance is expressed in a kinematic Taylor series to avoid making assumptions about the matter content or dynamics, and the inferred acceleration is essentially in the direction of the local peculiar bulk flow. Other works that argue for negligible evidence for anisotropy Brout:2022vxf ; Akarsu:2019pwn typically assume an isotropic framework to some degree or make corrections for peculiar velocities by transforming to the “CMB frame”, which removes the evidence for anisotropy. In Secrest:2020has , the authors show that the distribution of distant matter as traced by radio sources and quasars is not isotropic in the CMB frame. The authors of Secrest:2022uvx demonstrate a $5.1\sigma$ rejection of the null hypothesis that there is a dipole anisotropy in the distribution of radio galaxies and quasars consistent with our inferred peculiar motion. The question of bulk anisotropy may not be settled in the literature, but at the very least, metrics that account for anisotropy should be considered and studied.
The CMB constraint on anisotropy is very stringent, constraining any shift away from isotropy in a particular direction of expansion to be $\lesssim 10^{-7}$ Schucker:2014wca . The Planck satellite data supports at most a very tiny amount of anisotropy Planck:2019evm , and Hertzberg:2024uqy show that any anisotropy in the early universe is washed out by expansion. However, the constraints coming from the early-time CMB need not agree with late-time constraints from supernovae given the Hubble tension and growing support for MOND-like modified gravity models that can explain small scales well and may induce local anisotropies in structure formation, and since gravitational waves propagate through a background that is late-time, only the supernovae constraints were considered in this work. Another point to drive home is that the community is eager to estimate a distance-ladder-free estimate of the Hubble parameter from gravitational waves, but the inference of it may be incorrect due to the potential error in the inference of luminosity distance, which this work endeavors to explore.
@@
The accuracy of predictions of the waveforms of GWs due to massive binary systems has been refined to a high level over the years through advancements in numerical relativity. In the duration between emission of these waves from the source and detection by a gravitational wave (GW) observatory, the waveform is to a large degree unaffected by the spacetime through which it propagates. However, as more detections of GWs are made and more observatories come online with higher precision, our ability to discern subtle effects on the waveforms will increase. Many works have studied gravitational waves, including their exact solutions in some cases, in background spacetimes other than FLRW, such as Bianchi-type backgrounds 11 ; 12 ; 13 . Others have examined the effect of propagation through cosmic dust and in the presence of electromagnetic fields and scalar fields 21 ; 22 ; 23 , and much analysis on gravitational waves has been done in modified gravity models 31 ; 32 ; 33 . Our goal in this work is different in that we specifically want a solution that matches with the parameters of the source binary system without having to resort to an intensive numerical solution while also getting a solution easily comparable to the FLRW background case. We want this in order to make the point that one should be cautious in inferring luminosity distance while taking the usual FLRW background for granted.
In this work, using supernova anisotropy data as a reference, we show that the assumption of a Bianchi I background spacetime could imply an average of 2.1% difference (up to 5.9% depending on the direction of observation) in inferred luminosity distance compared to what would be inferred under the assumption of the FLRW background. The effect on the waveform is primarily in the amplitude as opposed to a more subtle effect on the oscillation pattern.
@@ II Bianchi I Metric
The simplest possible expanding universe that obeys the cosmological principle (homogeneous, isotropic) is described by the FLRW metric,
@@
where $a$, $b$ and $c$ are arbitrary functions of time. We have chosen to work with a perturbative form of this metric for the ease of comparing the waveform to the solution with an FLRW background, where $b$ and $c$ differ from $a$ only slightly:
@@
$c(t)=a(t)(1+\gamma),$
where $\beta\ll 1$ and $\gamma\ll 1$. In what follows, we keep to first order in $\beta$ and $\gamma$. In general, $\beta$ and $\gamma$ are functions of time, but we also assume here that $\beta$ and $\gamma$ are slowly varying: $\dot{\beta},\dot{\gamma}\ll 1$.
@@ II.1 Gravitational Waves
To obtain the wave equation, we solve Einstein’s equation to first order as outlined in Mondal:2020iet . The Riemann tensor is written as
$R^{\rho}_{\mu\nu\lambda}=\bar{R}^{\rho}_{\mu\nu\lambda}+R^{(1)\rho}_{\mu\nu\lambda},$ (5)
and the Einstein tensor is written as
$G_{\alpha\beta}=\bar{G}_{\alpha\beta}+G^{(1)}_{\alpha\beta}=8\pi(\bar{T}_{\alpha\beta}+T^{(1)}_{\alpha\beta}).$ (6)
Since the perturbation to the stress-energy tensor $T^{(1)}_{\alpha\beta}$ is primarily a long-wavelength correction, it is decoupled from the short-wavelength regime and irrelevant for the evolution of gravitational waves of observational interest, as discussed in Maggiore . We use the Lorenz gauge
$\nabla_{\mu}\tilde{h}^{\mu\nu}=0,$ (7)
where $\tilde{h}_{\mu\nu}\equiv h_{\mu\nu}-\frac{1}{2}\gamma_{\mu\nu}h$, and $h=h^{\alpha}_{\alpha}$, the trace of $h_{\mu\nu}$. In this case, Einstein’s equation at first order is $G^{(1)}_{\alpha\beta}=0$, given by Eq. (14) of Mondal:2020iet :
$\square\tilde{h}_{\beta\alpha}-2\bar{R}_{\delta\beta\alpha\rho}\tilde{h}^{\delta\rho}-\bar{R}_{\delta\alpha}\tilde{h}^{\delta}_{\beta}-\bar{R}_{\delta\beta}\tilde{h}^{\delta}_{\alpha}=0,$ (8)
where $\square\equiv\nabla^{\alpha}\nabla_{\alpha}$.
Without loss of generality, we consider a wave propagating in the $x$ direction. We also use the transverse-traceless (TT) gauge, in which $\tilde{h}_{\mu\nu}=h_{\mu\nu}$. The propagation in the $x$ direction, the transverse condition, the traceless condition, and the symmetry of the metric tensor imply the following conditions:
@@
Then $h_{\mu\nu}$ is
$\begin{pmatrix}0&0&0&0\\ 0&0&0&0\\ 0&0&h_{22}&h_{23}\\ 0&0&-h_{23}&h_{22}.\end{pmatrix}$ (10)
As a result, Eq. (8) gives two non-zero independent equations. Using Eq. (8), $(G^{(1)}_{22}-G^{(1)}_{33})/2$=0 gives
$-\left(\frac{\dot{b}^{2}}{b^{2}}+\frac{b\dot{b}\dot{c}}{c^{3}}+\frac{c\dot{b}\dot{c}}{b^{3}}+\frac{\dot{c}^{2}}{c^{2}}\right)h_{22}+\left(\frac{1}{a^{2}}\frac{\partial^{2}}{\partial x^{2}}+\frac{1}{b^{2}}\frac{\partial^{2}}{\partial y^{2}}+\frac{1}{c^{2}}\frac{\partial^{2}}{\partial z^{2}}\right)h_{22}+\left(-\frac{\dot{a}}{a}+\frac{\dot{b}}{b}+\frac{\dot{c}}{c}\right)\dot{h}_{22}-\ddot{h}_{22}=0,$ (11)
where $\dot{}$ denotes differentiation with respect to $t$, and $(G^{(1)}_{23}+G^{(1)}_{32})/2$=0 gives
$-\frac{4\dot{b}\dot{c}}{bc}h_{23}+\left(\frac{1}{a^{2}}\frac{\partial^{2}}{\partial x^{2}}+\frac{1}{b^{2}}\frac{\partial^{2}}{\partial y^{2}}+\frac{1}{c^{2}}\frac{\partial^{2}}{\partial z^{2}}\right)h_{23}+\left(-\frac{\dot{a}}{a}+\frac{\dot{b}}{b}+\frac{\dot{c}}{c}\right)\dot{h}_{23}-\ddot{h}_{23}=0.$ (12)
Under the “local wave approximation” (which allowed us to use $G^{(1)}_{\mu\nu}=0$), the background dynamics are slow compared to the angular frequency $\omega$ of the wave:
$\mathcal{O}\left(\frac{1}{t}h_{\mu\nu}\right)\ll\mathcal{O}(\omega h_{\mu\nu}).$ (13)
The time derivative of any of the scale factors goes like $1/t$, so the first and third sets of terms of Eqs. (11) and (12) go like $h_{\mu\nu}/t^{2}$ and $\omega h_{\mu\nu}/t$ respectively and are negligible compared to the other terms, which are second derivatives of $h_{\mu\nu}$ and go like $\omega^{2}h_{\mu\nu}$. Therefore, Eqs. (11) and (12) simplify to the same form:
$\frac{1}{a^{2}}\left(\frac{d^{2}}{dx^{2}}+\frac{1}{(1+\beta)^{2}}\frac{d^{2}}{dy^{2}}+\frac{1}{(1+\gamma)^{2}}\frac{d^{2}}{dz^{2}}\right)\phi-\ddot{\phi}=0,$ (14)
where $\phi(\eta,x,y,z)$ stands in for $h_{22}(\eta,x,y,z)$ and $h_{32}(\eta,x,y,z)$.
@@
$\frac{1}{a^{2}}\left(\frac{d^{2}}{dx^{2}}+\frac{1}{(1+\beta)^{2}}\frac{d^{2}}{dy^{2}}+\frac{1}{(1+\gamma)^{2}}\frac{d^{2}}{dz^{2}}\right)\phi+\frac{a^{\prime}}{a^{3}}\phi^{\prime}-\frac{1}{a^{2}}\phi^{\prime\prime}=0,$ (16)
where denotes differentiation with respect to $\eta$. We write this equation as
$\frac{1}{a^{2}}\tilde{\nabla}^{2}\phi+\frac{a^{\prime}}{a^{3}}\phi^{\prime}-\frac{1}{a^{2}}\phi^{\prime\prime}=0,$ (17)
@@
~$\tilde{\nabla}^{2}\equiv\frac{d^{2}}{dx^{2}}+\frac{1}{(1+\beta)^{2}}\frac{d^{2}}{dy^{2}}+\frac{1}{(1+\gamma)^{2}}\frac{d^{2}}{dz^{2}}.$ (18)
The second term in Eq. (17) goes like $\omega h_{\mu\nu}/t$ and, as before, is negligible. Therefore, the equation to solve is now
~$\tilde{\nabla}^{2}\phi-\phi^{\prime\prime}=0.$ (19)
@@
$\phi(\eta,x,y,z)=\frac{g(\eta-\tilde{r})}{\tilde{r}a(\eta)},$ (20)
@@
This $\tilde{r}$ is very similar to the Cartesian $r=\sqrt{x^{2}+y^{2}+z^{2}}$ in the FLRW background. The factors of $\beta$ and $\gamma$ which scale the scaling factor $a$ for the $y$ and $z$ directions can be thought of equivalently as rescalings of $y$ and $z$:
@@
~$\tilde{\nabla}^{2}\left(\frac{g}{\tilde{r}a}\right)-\left(\frac{g^{\prime\prime}}{\tilde{r}a}-\frac{g\tilde{r}^{\prime\prime}}{\tilde{r}^{2}a}+2\frac{g\tilde{r}^{\prime 2}}{\tilde{r}^{3}a}+\frac{g\tilde{r}^{\prime}a^{\prime}}{\tilde{r}^{2}a^{2}}-\frac{ga^{\prime\prime}}{\tilde{r}a^{2}}+\frac{ga^{\prime}\tilde{r}^{\prime}}{\tilde{r}^{2}a^{2}}+2\frac{ga^{\prime 2}}{\tilde{r}a^{3}}\right)=0.$ (23)
From Eq. (21), we see that $\tilde{r}^{\prime}$ is non-zero only at or above the first-order level in $\dot{\beta}$ and $\dot{\gamma}$. Within the second set of parentheses of the above equation, all terms go like $h_{\mu\nu}/t^{2}$ except for the first term, so we keep only that first term. We now have the final form of the wave equation to solve:
~$\tilde{\nabla}^{2}\left(\frac{g}{\tilde{r}a}\right)-\frac{g^{\prime\prime}}{\tilde{r}a}=0.$ (24)
The left-hand side of Eq. (24), to first order in $\beta$, $\gamma$, $\dot{\beta}$, and $\dot{\gamma}$, is
$\frac{2y^{2}\beta^{\prime}g^{\prime\prime}(\eta-r)}{r^{2}a}+\frac{2z^{2}\gamma^{\prime}g^{\prime\prime}(\eta-r)}{r^{2}a}-\frac{y^{2}\beta^{\prime\prime}g^{\prime}(\eta-r)}{r^{2}a}-\frac{z^{2}\gamma^{\prime\prime}g^{\prime}(\eta-r)}{r^{2}a}.$ (25)
Given that $g=\tilde{r}a\phi$ (from Eq. (20)), we see that $g^{\prime}\sim\omega g$ for its biggest contributing term, and each term in Eq. (25) is thus negligible under the local wave approximation since $\beta^{\prime},~{}\gamma^{\prime}\sim\beta/t,~{}\gamma/t$. Therefore, the ansatz Eq. (20) satisfies Eq. (24). The solution for the corresponding GW equation that results when an FLRW background is used (see Eq. 4.182 of Maggiore ) is the same as the solution Eq. (20) when
@@
The solution for a generic binary system (black hole-black hole, neutron star-black hole, neutron star-neutron star, etc) in circular orbit to first order in $v/c$ can be made to include back-reaction with respect to an FLRW background in a simple way using the local wave approximation (Eqs. 4.191 and 4.192 from Maggiore ). Because we have the same solution if we substitute $\tilde{r}$ for $r$, we can also express the solution for the two polarizations of GWs for our perturbative Bianchi I (PBI) background in a similar manner:
$h_{PBI,+}=\frac{GM_{c}(1+\mathrm{z})/c^{2}}{a(t_{0})\tilde{r}(1+\mathrm{z})}\left(\frac{5GM_{c}(1+\mathrm{z})/c^{3}}{t_{coal}-t}\right)^{1/4}\cos{\left(-2\left(\frac{t_{coal}-t}{5GM_{c}(1+\mathrm{z})/c^{3}}\right)^{5/8}+\Phi_{0}\right)}\frac{1+\cos^{2}\iota}{2},$ (27)
$h_{PBI,\times}=\frac{GM_{c}(1+\mathrm{z})/c^{2}}{a(t_{0})\tilde{r}(1+\mathrm{z})}\left(\frac{5GM_{c}(1+\mathrm{z})/c^{3}}{t_{coal}-t}\right)^{1/4}\sin{\left(-2\left(\frac{t_{coal}-t}{5GM_{c}(1+\mathrm{z})/c^{3}}\right)^{5/8}+\Phi_{0}\right)}\cos\iota,$ (28)
where $t_{0}$ is the present time, $M_{c}$ is the chirp mass (a combination of the reduced mass and the total mass of the binary system), $\mathrm{z}$ is the redshift of the system, $t_{coal}$ is the coalescence time, $\Phi_{0}$ is an integration constant specifying the phase of the wave, and $\iota$ is the angle of inclination of the orbital plane of the system. Only the first factor of these equations will be relevant for this work as the luminosity distance only depends on the first factor. The cosine function containing $\eta_{obs}$ is the oscillatory part, while the the other factors determines the evolution of the envelope of the oscillation. The luminosity distance in FLRW space is
$d_{FLRW}(\mathrm{z})=a(t_{0})r(1+\mathrm{z}),$ (29)
@@
$h_{FLRW,+}=\frac{GM_{c}(1+\mathrm{z})/c^{2}}{d_{FLRW}(\mathrm{z})}\left(\frac{5GM_{c}(1+\mathrm{z})/c^{3}}{t_{coal}-t}\right)^{1/4}\cos{\left(-2\left(\frac{t_{coal}-t}{5GM_{c}(1+\mathrm{z})/c^{3}}\right)^{5/8}+\Phi_{0}\right)}\frac{1+\cos^{2}\iota}{2},$ (30)
$h_{FLRW,\times}=\frac{GM_{c}(1+\mathrm{z})/c^{2}}{d_{FLRW}(\mathrm{z})}\left(\frac{5GM_{c}(1+\mathrm{z})/c^{3}}{t_{coal}-t}\right)^{1/4}\sin{\left(-2\left(\frac{t_{coal}-t}{5GM_{c}(1+\mathrm{z})/c^{3}}\right)^{5/8}+\Phi_{0}\right)}\cos\iota.$ (31)